OpenStudy (brooke2353):

Three brothers live in a farm. They agreed to buy new seeds: Adam and Ben would go and Charlie stayed to protect fields. Ben bought 75 sacks of wheat in the market whereas Adam bought 45 sacks. At home, they split the sacks equally. Charlie had paid 1400 dollars for the wheat. How much dollars did Ben and Adam get of the sum, considering equal split of the sacks?

OpenStudy (brooke2353):

@dm91

OpenStudy (brooke2353):

@DCchamp03

OpenStudy (brooke2353):

@hysusonic

OpenStudy (dm91):

what do you think it is?

OpenStudy (brooke2353):

UGHH i think Ben spent $1225 and Adam spent $175

OpenStudy (dm91):

okay you have $1400 that the spent total

OpenStudy (dm91):

Ben spent the most because he got 75 sacks

OpenStudy (brooke2353):

so then Every farmer's part is 1/3(45+75) = 40 sacks. Charlie paid $1400 for 40 sacks, then 1 sack costs $1400/40 = $35/sack. Adam got $35*(45-40)=35*5 = $175. Ben got $35*(75-40)=35*35 = $1225

OpenStudy (dm91):

so yeah that would be the right answer that you are thinking

OpenStudy (brooke2353):

okay

OpenStudy (dm91):

yeah that would be it right there brooke

OpenStudy (dm91):

if you need help with anything just tag me in it and i will try to help